# SplatFormer: Point Transformer for Robust 3D Gaussian Splatting

### v2

@@ 5 Experiments
Limitations and Future Work. Our method has several limitations that provide directions for future work. First, despite outperforming all the considered baselines, it still struggles to reconstruct fine-grained details and complex texture. Second, the generalization to real-world captures could be improved by scaling up training examples and by enhancing the realism of synthetic lighting. Third, applying our method to refining 2DGS may further improve the OOD-NVS results. Finally, it would be valuable to train our method to remove OOD-NVS artifacts in unbounded scenes and with a wider range of OOD camera setups. In the appendix F, we present a experimental result in Fig. F.1 and Tab. F.1 using the MVImgNet dataset  (Yu et al., 2023), and outline both the potential and challenges. Please refer to it for an extended discussion.
@@ 6 Conclusion
Funding
This study was conducted within the national ”Proficiency” [^1] https://surgicalproficiency.ch research project funded by the Swiss Innovation Agency Innosuisse in 2021 as one of 15 flagship initiatives.
@@
We provide further details on the evaluation datasets (Sec. A), implementation specifics (Sec. B), and ablation experiments (Sec. C). Additionally, we describe baseline implementations (Sec. D) and present comprehensive visual comparisons (Sec. E). Finally, we discuss the limitations of SplatFormer in Sec. F.
@@ Appendix A Evaluation Datasets
Synthetic Datasets. We use Blender to render objects from ShapeNet (Chang et al., 2015), Objaverse-v1 (Deitke et al., 2023), and GSO (Google Scanned Objects) (Downs et al., 2022). The camera setups for the three evaluation sets are consistent: $N_{in}=32$ input views cover the $360\degree$ azimuth and elevation angle $\phi$ varies according to a sinusoidal function ranging between $(0,\phi_{\textrm{max}})$. For each object in ShapeNet, we rotate the object’s shortest side with the z-axis and render a single set of input views with $\phi_{\textrm{max}}=10\degree$. For each object in Objaverse-v1 and GSO, we render two sets of input views with $\phi_{\textrm{max}}=10\degree,20\degree$. The out-of-distribution (OOD) test set consists of $N_{\textrm{out}}=9$ views, with $\phi_{\textrm{ood}}=(70\degree,80\degree,90\degree)$ and randomly sampled azimuths. The rendered resolution is $256\times 256$ pixels. The resulting ShapeNet-OOD, Objaverse-OOD, GSO-OOD datasets include a total of 20, 40, and 40 input-test experiments, respectively. We enable specular effects to achieve more realistic rendering results when using objects from Objaverse-v1 and GSO, and disable specular reflections for ShapeNet to study a more basic illumination setup.
Real-world OOD iPhone Dataset. We have captured 4 scenes featuring an object of interest using an iPhone, with the images and camera setups shown in Fig. F.5. Each scene contains around 30 input views and 4 OOD test views. During evaluation, we first generate foreground masks of the objects of interest for the OOD test view using SAM2 (Ravi et al., 2024), and then only evaluate the pixels within the mask. To refine the 3DGS representation via SplatFormer, we crop out the part of the 3DGS point cloud that corresponds to the foreground region using selection tools in MeshLab. This may also be easily done by automatic 3D detection methods like Segment3D (Huang et al., 2024b). The cropped splats are then refined via SplatFormer and rendered using the standard 3DGS-based rendering pipeline. We resize images to the resolution of $300\times 400$ for both 3DGS training and evaluation.
@@ Appendix B Implementation Details
Network Arhitecture. The point transformer encoder begins with an MLP embedding layer, followed by five down-pooling and four up-pooling stages, ultimately producing features with a dimensionality of $V=96$. The down-pooling stages contain $(2,2,2,6,2)$ attention blocks and have hidden dimensions of $(64,96,128,256,512)$. Each stage, except the first, is followed by a down-sampling grid-pooling layer. The up-pooling stages consist of $(2,2,2,2)$ attention blocks, with hidden dimensions of $(256,128,96,96)$. Each stage, except the first, is preceded by an up-sampling grid-pooling layer. A grid resolution of 384 is used to voxelize the point cloud, and the strides for the grid-pooling layers are set to $(1,2,2,2)$. For the architecture details of attention blocks and grid pooling please refer to Wu et al. (2024b).
@@
Training Dataset Curation. The ShapeNet training set contains 33k objects, all available for non-commercial research and educational purposes. The Objaverse training set includes 48k objects from Objaverse-1.0 (Deitke et al., 2023), all licensed under Creative Commons for distribution. We use Blender to render each object with 32 low-elevation views and 5 top-down views. Diffuse lighting and materials are applied in ShapeNet scenes, while specular effects and shadows are enabled in Objaverse scenes. For the rendered 2D low-elevation views, we use gsplat (Ye et al., 2024) to optimize the initial 3D Gaussian splats (3DGS) for each scene. The spherical harmonics degree is set to 0 for ShapeNet and 3 for Objaverse. To reduce computational costs, we terminate the optimization early at 10k steps, where evaluation performance levels off. We process the scenes using 48 RTX-2080Ti GPUs, with rendering and 3DGS optimization taking approximately 3 minutes per scene. It takes 2 days to generate each training dataset.
Training. For each scene, we render 4 target images at each iteration, with 70% OOD views and 30% input views, for photometric supervision. For the training of our full model, we use 8 RTX-3090s with one scene per GPU, set gradient accumulation steps as 4, and train for 150k iterations, which takes around 2 days. We use Adam optimizer with a constant learning rate of 3e-5.
Appendix C More Experiments
SplatFormer vs 2D Denoising. In addition to the metrics in Tab. 5, Fig. C.1 presents a visual comparison between the 2D denoising method DiffBIR (Lin et al., 2024) and SplatFormer on ShapeNet-OOD test views. Though the DiffBIR-stage1 model removes certain artifacts, the improvements are inconsistent across views. Retraining a 3DGS model using the generated images fails to fully address these limitations. Additionally, the stage-1 model struggles to infer correct geometry from the noisy 2D images, causing errors that propagate to stage-2, which may introduce unfaithful hallucinations. In contrast, our method processes both input and output in 3D, resulting in more accurate and consistent artifact removal.
@@ Appendix C More Experiments
Figure C.1: We adopt DiffBIR (Lin et al., 2024) to denoise artifacts in 2D space. Additionally, we retrain 3DGS using the denoised images to improve multi-view consistency. However, 2D denoising alone is insufficient for fully recovering geometry, as it relies solely on 2D inputs.
@@
Appendix D Baseline Implementations
MipNeRF360. We use the officially released code [^3] https://github.com/google-research/multinerf of MipNeRF360 (Müller et al., 2022) to produce its results on our OOD benchmarks. The model is trained using batch size 512 for 250k iterations, with learning rate 0.00025. The training of each scene takes approximately 7 hours on a single RTX-3090 GPU. The near and far planes used in volumetric rendering are determined by the ray bounding box intersection following the approach used in InstantNGP. The random background trick used in InstantNGP is also applied to help removing the floaters.
3D Gaussian Splatting. We use the gsplat (Ye et al., 2024) and the default hyperparameters of the toolbox. Specifically, we set the spherical harmonics degree to 3, the number of iterations to 30k, warm-up steps to 500, densifying and culling Gaussians every 500 step, and we stop the density control at 15k steps. For synthetic datasets, we use the ground truth camera poses provided by the Blender rendering. Given that COLMAP struggles to reconstruct certain synthetic objects due to symmetry and smooth textures, we follow the original 3D Gaussian Splatting (3DGS) approach (Kerbl et al., 2023) by randomly sampling 50,000 points within the bounding box of the objects. For our real-world iPhone captures, we first estimate camera poses using all views and then perform point triangulation, with only the input views used to estimate the initial point cloud.
InstantSplat (DUSt3R-inititialized 3DGS). InstantSplat (Fan et al., 2024) utilizes an end-to-end dense stereo model, namely DUSt3R (Wang et al., 2024), to estimate dense point clouds and camera poses for 3DGS initialization. To assess the effectiveness of dense initialization on our OOD-NVS task, we feed all input views, and the given camera poses into the DUSt3R reconstruction pipeline. Given the large number of training views, DUSt3R produces approximately 10 million points, which can lead to out-of-memory issues during 3DGS optimization. We randomly sample 100k points from the total point cloud as initialization for 3DGS. We avoid selecting the 100k points with the highest confidence as this would disproportionately retain floors and walls, which tend to have overly high confidence scores in DUSt3R. We then replace the random initialization in our 3DGS implementation with the dense point clouds, leaving the rest of the 3DGS optimization process unchanged.
Nerfbusters. We use the official codebase [^8] https://github.com/ethanweber/nerfbusters of Nerfbusters (Warburg* et al., 2023). We first train Nerfacto (Tancik et al., 2023) for 30k steps, and then run Nerfbusters on pretrained Nerfacto models for 5k steps to remove the artifacts learned by Nerfacto. The same near and far planes calculation strategy described in MipNeRF360 is employed.
SyncDreamer. We use the official codebase [^9] https://github.com/liuyuan-pal/SyncDreamer of SyncDreamer (Liu et al., 2023b). SyncDreamer only supports single input view, which is far from sufficient to reconstruct the feature volume needed for the out-of-distribution view synthesis. Since the focus point of our training cameras are not located at the origins, we remove the camera embedding vectors which rely on relative azimuths and elevations in their synthetic setup. Note that training SyncDreamer requires significantly greater computational resources (multiple Nvidia A100s), which is not comparable with ones required to train our method and other baselines (RTX-3090s).
LaRa. We use LaRa’s released codebase [^10] https://github.com/autonomousvision/LaRa and train two models on the same ShapeNet-OOD and Objaverse-OOD training sets as our method and evaluate their performances on the respective benchmarks. Due to hardware limitations, LaRa can only process a maximum of four input views on RTX-3090 GPUs. To maximize scene coverage, we randomly select the four most widely separated input views. For training, we uniformly sample from all available views, incorporating both in-distribution and out-of-distribution views for the 2D image loss. During inference, we feed the model four input views and rasterize the predicted Gaussian primitives to the OOD test views. While it is possible to divide the input views into groups of four, and then combine the processed resulting primitives for rasterization, this approach compromises global consistency and yields worse performance compared to using just four views directly. Following LaRa’s default setup (Chen et al., 2024a), we train each model for 150k iterations using 8 RTX 3090 GPUs, which takes approximately 2 days. We initialize training with LaRa’s pretrained weights, which speeds up convergence and improves performance.
Other Baselines. We acknowledge that some related approaches are not included in this paper due to redundancy or infeasibility. DNGaussian (Li et al., 2024a) supervises 3DGS training with depth maps predicted by monocular depth estimators. The idea is similar to its concurrent work FSGS (Zhu et al., 2024) which, in our OOD-NVS dense capture setup, does not outperform 3DGS. CAT3D (Gao et al., 2024) and ReconFusion (Wu et al., 2024a) use internal image diffusion models for initialization but neither has released code and models. GeNVS (Chan et al., 2023) and 3DiM (Watson et al., 2023) are also publicly unavailable. To compare with methods that utilize pretrained 2D diffusion models, we evaluate SyncDreamer (Liu et al., 2023b) which is designed to achieve better multi-view consistency. Other similar approaches like ZeroNVS (Sargent et al., 2024), Zero123 (Liu et al., 2023a), and Vivid123 (Kwak et al., 2024), only support single-view input. We also tried ViewDiff (Höllein et al., 2024), a similar 2D-prior method where each model is trained for a single object category. However, we found that ViewDiff could not converge when trained on multiple object categories in our cases. Lastly, in the category of learning-based 2D-to-3D methods, SpaRP (Xu et al., 2024), MVSplat (Chen et al., 2024b), and LatentSplat (Wewer et al., 2024) are constrained by the memory limitations of RTX 3090 GPUs, allowing a maximum of four input views. We found that training these methods on our dataset with four large-baseline input views, as we did with LaRa, resulted in failure. This is because these methods rely on overlapping input views to compute cross-view correspondences. We also attempted to partition the consecutive input views into groups and aggregate the predicted Gaussian splats after running each group independently. However, this approach introduced significant global inconsistencies. Consequently, we do not report their results due to their incompatibility with our OOD-NVS task.
Appendix E Comparisons with Baselines
@@
Evaluation Details. As mentioned in Sec. A, we create two experimental setups for each scene in Objaverse and GSO, with the same test views yet different input views with maximum elevations $\phi_{\textrm{max}}=10\degree$ and $20\degree$ respectively. The final evaluation scores reported in Tab. 1 and Tab. 5 are averaged across the two sets of experiments. We also report the separate scores of the two sets in Tab. E.1 (Objaverse-OOD) and Tab. E.2 (GSO-OOD). The evaluation scores of $\phi_{\textrm{max}}=20\degree$ are better than $\phi_{\textrm{max}}=10\degree$ as the input viewing angles are slightly closer to the OOD test views. However in both setups, the elevation deviation, with $\phi_{\textrm{ood}}\geq 70\degree$, is quite large and our method outperforms the baselines consistently.
@@
Table E.1: Detailed Results on Objaverse-OOD. We report the separate OOD evaluation results of the two sets of experiments with input views’ maximum elevations $\phi_{\textrm{max}}=(10\degree,20\degree)$. The average results are used in Tab. 1. Colors indicate the 1st, 2nd, and 3rd best-performing models.
@@
Nerfbusters (Warburg* et al., 2023) 15.25 0.643 0.318 18.49 0.735 0.255 16.87 0.689 0.287
SyncDreamer (Liu et al., 2023b) 9.64 0.423 0.487 9.81 0.418 0.491 9.72 0.421 0.489
@@
Table E.2: Detailed Results on GSO-OOD. We report the separate OOD evaluation results of the two sets of experiments with input views’ maximum elevations $\phi_{\textrm{max}}=(10\degree,20\degree)$. The average results are used in Tab. 5. Colors indicate the 1st, 2nd best-performing models.
@@
Visual Comparisons. We show more visual comparisons with all the baselines in Fig. F.2 (ShapeNet-OOD), Fig. F.3 (Objaverse-OOD), Fig. F.4 (GSO-OOD), and Fig. F.5 (Real-world OOD). We also provide a supplementary video to show the comparisons.
@@
For prior-enhanced NVS methods, SSDNeRF (Chen et al., 2023), which is designed to learn category-specific object priors in its original paper, struggles to learn cross-category priors in our training set, resulting in severe artifacts. Nerfbusters (Warburg* et al., 2023) incorrectly identifies many structures as floaters, leading to the removal of significant parts of objects. While FSGS (Zhu et al., 2024) achieves a more balanced densification of Gaussians and smoother depth regularization, it only offers marginal visual improvements over 3DGS and continually fails in certain scenarios. InstantSplat (Fan et al., 2024) also cannot mitigate the artifacts caused by overfitting by using a dense point cloud initialization.
@@
Among learning-based feed-forward methods, LaRa (Chen et al., 2024a) can infer plausible geometry and textures from input views. However, its reliance on a limited number of input views and voxel representation restricts its ability to process and represent high-frequency details. SyncDreamer (Liu et al., 2023b) fails to synthesize the deviated test view as it only supports single input view which is quite insufficient for OOD-NVS.

### v3

@@ 5 Experiments
Limitations and Future Work. Our method has several limitations that provide directions for future work. First, despite outperforming all the considered baselines, it still struggles to reconstruct fine-grained details and complex texture. Second, the generalization to real-world captures could be improved by scaling up training examples and by enhancing the realism of synthetic lighting. Third, applying our method to refining 2DGS may further improve the OOD-NVS results. Finally, it would be valuable to train our method to remove OOD-NVS artifacts in unbounded scenes and with a wider range of OOD camera setups. In the appendix H, we present a experimental result in Fig. H.1 and Tab. H.1 using the MVImgNet dataset  (Yu et al., 2023), and outline both the potential and challenges. Please refer to it for an extended discussion.
@@ 6 Conclusion
Acknowledgements
This study was conducted within the national “Proficiency” [^1] https://surgicalproficiency.ch research project funded by the Swiss Innovation Agency Innosuisse in 2021 as one of 15 flagship initiatives. This work was also supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a03 on Alps. Marko Mihajlovic is in part supported by the Hasler Stiftung Grant (2024-09-12-159).
@@ Acknowledgements
We provide details on the evaluation datasets (Sec. A) and implementations of our method (Sec. B). Then, we show more experimental results, including ablation studies (Sec. C), evaluation on geometry (Sec. D), evaluation on a diverse range of test views (Sec. D) and comprehensive visual comparisons (Sec. F). Additionally, we describe baseline implementations (Sec. G). Finally, we discuss the limitations of SplatFormer in Sec. H.
@@ Appendix A Evaluation Datasets
Synthetic Datasets. We use Blender to render objects from ShapeNet (Chang et al., 2015), Objaverse-v1 (Deitke et al., 2023), and GSO (Google Scanned Objects) (Downs et al., 2022). The camera setups for the three evaluation sets are consistent: $N_{\textrm{in}}=32$ input views cover the $360\degree$ azimuth and elevation angle $\phi$ varies according to a sinusoidal function ranging between $(0,\phi_{\textrm{max}})$. For each object in ShapeNet, we rotate the object’s shortest side with the z-axis and render a single set of input views with $\phi_{\textrm{max}}=10\degree$. For each object in Objaverse-v1 and GSO, we render two sets of input views with $\phi_{\textrm{max}}=10\degree,20\degree$. The out-of-distribution (OOD) test set consists of $N_{\textrm{out}}=9$ views, with $\phi_{\textrm{ood}}=(70\degree,80\degree,90\degree)$ and uniformly strided azimuths. The rendered resolution is $256\times 256$ pixels. The resulting ShapeNet-OOD, Objaverse-OOD, GSO-OOD datasets include a total of 20, 40, and 40 input-test experiments, respectively. We enable specular effects to achieve more realistic rendering results when using objects from Objaverse-v1 and GSO, and disable specular reflections for ShapeNet to study a more basic illumination setup.
Real-world OOD iPhone Dataset. We have captured 4 scenes featuring an object of interest using an iPhone, with the images and camera setups shown in Fig. H.6. Each scene contains around 30 input views and 4 OOD test views. During evaluation, we first generate foreground masks of the objects of interest for the OOD test view using SAM2 (Ravi et al., 2024), and then only evaluate the pixels within the mask. To refine the 3DGS representation via SplatFormer, we crop out the part of the 3DGS point cloud that corresponds to the foreground region using selection tools in MeshLab. This may also be easily done by automatic 3D detection methods like Segment3D (Huang et al., 2024b). The cropped splats are then refined via SplatFormer and rendered using the standard 3DGS-based rendering pipeline. We resize images to the resolution of $300\times 400$ for both 3DGS training and evaluation.
@@ Appendix B Implementation Details
Network Arhitecture. The point transformer encoder begins with an MLP embedding layer, followed by five down-pooling and four up-pooling stages, ultimately producing features with a dimensionality of $V=96$. The down-pooling stages contain $(2,2,2,6,2)$ attention blocks and have hidden dimensions of $(64,96,128,256,512)$. Each down-pooling stage, except the first, is followed by a down-sampling grid-pooling layer. The up-pooling stages consist of $(2,2,2,2)$ attention blocks, with hidden dimensions of $(256,128,96,96)$. Each up-pooling stage, except the last, is preceded by an up-sampling grid-pooling layer. A grid resolution of 384 is used to voxelize the point cloud, and the strides for the grid-pooling layers are set to $(1,2,2,2)$. For the architecture details of attention blocks and grid pooling please refer to Wu et al. (2024b).
@@
Training Dataset Curation. The ShapeNet training set contains 33k objects, all available for non-commercial research and educational purposes. The Objaverse training set includes 48k objects from Objaverse-1.0 (Deitke et al., 2023), all licensed under Creative Commons for distribution. We use Blender to render each object with 32 low-elevation views and 5 top-down views. Diffuse lighting and materials are applied in ShapeNet scenes, while specular effects and shadows are enabled in Objaverse scenes. For the rendered 2D low-elevation views, we use gsplat (Ye et al., 2024) to optimize the initial 3D Gaussian splats (3DGS) for each scene. The spherical harmonics degree is set to 0 for ShapeNet and 1 for Objaverse. To reduce computational costs, we terminate the optimization early at 10k steps, where evaluation performance levels off. We process the scenes using 48 RTX-2080Ti GPUs, with rendering and 3DGS optimization taking approximately 3 minutes per scene. It takes 2 days to generate each training dataset.
Training. For each scene, we render 4 target images at each iteration, with 70% OOD views and 30% input views, for photometric supervision. For the training of our full model, we use 8 RTX-4090s with one scene per GPU, set gradient accumulation steps as 4, and train for 150k iterations, which takes around 2 days. We use Adam optimizer with a constant learning rate of 3e-5. During training, we cap the number of input Gaussians to SplatFormer at 100k. If the 3DGS exceeds this threshold, we randomly subsample the Gaussians.
Inference. To obtain the refined 3DGS for each test scene, we first train a 3DGS with 32 input views using gsplat (Ye et al., 2024) for 10k steps. Then, we feed the 3DGS into SplatFormer to obtain the refined output. Regarding SplatFormer’s inference efficiency, most input splats in our object-centric test sets contain 70k—100k gaussians, requiring only 900MB of GPU memory for one feed-forward inference pass and achieving an inference time of 108 ms. To evaluate the upper bound of SplatFormer’s inference capability, we increase the number of Gaussians by sampling additional Gaussians with Gaussian noise. We find that an RTX 4090 GPU can accommodate up to 4 million Gaussians. However, it is important to note that the GPU memory consumption of the point transformer is not solely determined by the number of input points. Instead, it is also significantly influenced by the spatial distribution of the points. A 3DGS with a spatially uniform distribution and high entropy tends to consume more GPU memory than one with a more concentrated distribution. Since object-centric scenes often possess concentrated spatial distribution, our current SplatFormer can be quite efficient for large-scale 3DGS during inference. The primary computational bottleneck still lies in the training stage. Further improving the efficiency of point transformer for large-scale unbounded scenes remains an important direction for future work.
Appendix C Additional Ablation Studies
SplatFormer vs 2D Denoising. In addition to the metrics in Tab. 5, Fig. C.1 presents a visual comparison between the 2D denoising method DiffBIR (Lin et al., 2021) and SplatFormer on ShapeNet-OOD test views. Though the DiffBIR-stage1 model removes certain artifacts, the improvements are inconsistent across views. Retraining a 3DGS model using the generated images fails to fully address these limitations. Additionally, the stage-1 model struggles to infer correct geometry from the noisy 2D images, causing errors that propagate to stage-2, which may introduce unfaithful hallucinations. In contrast, our method processes both input and output in 3D, resulting in more accurate and consistent artifact removal.
@@ Appendix C Additional Ablation Studies
Figure C.1: We adopt DiffBIR (Lin et al., 2021) to denoise artifacts in 2D space. Additionally, we retrain 3DGS using the denoised images to improve multi-view consistency. However, 2D denoising alone is insufficient for fully recovering geometry, as it relies solely on 2D inputs.
@@
Appendix D Geometry Results
Table D.1: Geometry Evaluation on Objaverse-OOD.
[…]
3DGS 6.70e-4 0.239
Ours 4.05e-4 0.214
Building upon 3DGS, SplatFormer focuses primarily on enhancing novel view synthesis rather than surface extraction. However, we demonstrate that our method can still refine the geometry of input 3D Gaussians. Tab. D.1 compares the mean absolute errors (MAE) of rendered depths and normals under out-of-distribution (OOD) views between 3DGS and our approach. Specifically, depth maps for both methods are obtained as the weighted average depth of Gaussian primitives, following the standard approach implemented in the gsplat toolbox (Ye et al., 2024). Normal maps are then computed using finite differences on the estimated surface derived from the depth maps, as in 2DGS (Huang et al., 2024a). Fig. H.3 visualizes the depth and normal estimations, highlighting both quantitative and qualitative improvements achieved by our method over 3DGS.
Appendix E Evaluation Across Diverse Test Views
Though Tab. 1 and Tab. 5 only evaluate views with elevation $\phi_{\textrm{ood}}=(70\degree,80\degree,90\degree)$ and camera-to-origin distance $R=1$, SplatFormer can also enhance a wide range of test views with various elevations and even extreme close-up views. Fig. 2 shows that SplatFormer consistently outperforms 3DGS at elevation angles between $20\degree$ and $90\degree$. To further demonstrate this, we compare the PSNRs for views with different elevations $\phi\in[20\degree,90\degree]$ and camera radii ($R\in[0.2,1.0]$) between 3DGS and SplatFormer on the GSO-OOD dataset. As shown in Tab. E.1, SplatFormer significantly outperforms 3DGS across various viewing angles and even in extreme close-up views ($R=0.2$).
Table E.1: Results on various test views. We evaluate the PSNRs on novel views with various elevation angles $\phi$ and camera-to-origin distance $R$ in GSO-OOD test sets. Our method, SplatFormer, outperforms 3DGS (Kerbl et al., 2023) across various viewing angles and even in zoomed-in views.
$\phi=20\degree$ $\phi=30\degree$ $\phi=40\degree$ $\phi=50\degree$ $\phi=60\degree$ $\phi=70\degree$ $\phi=80\degree$ $\phi=90\degree$
$R=0.2$ 3DGS 22.29 21.13 20.73 19.17 18.60 18.00 17.46 16.84
Ours 22.99 22.19 21.85 21.50 21.44 21.18 20.89 20.18
[…]
Ours 23.74 23.14 22.44 21.76 21.38 21.23 21.00 21.04
[…]
Ours 25.33 24.04 23.19 22.71 22.43 22.24 22.14 22.22
[…]
Ours 27.28 25.87 24.85 24.21 23.89 23.68 23.57 23.61
[…]
Ours 28.97 27.62 26.54 25.77 25.34 25.08 24.93 25.03
Appendix F Comparisons with Baselines
@@
Evaluation Details. As mentioned in Sec. A, we create two experimental setups for each scene in Objaverse and GSO, with the same test views yet different input views with maximum elevations $\phi_{\textrm{max}}=10\degree$ and $20\degree$ respectively. The final evaluation scores reported in Tab. 1 and Tab. 5 are averaged across the two sets of experiments. We also report the separate scores of the two sets in Tab. F.1 (Objaverse-OOD) and Tab. F.2 (GSO-OOD). The evaluation scores of $\phi_{\textrm{max}}=20\degree$ are better than $\phi_{\textrm{max}}=10\degree$ as the input viewing angles are slightly closer to the OOD test views. However in both setups, the elevation deviation, with $\phi_{\textrm{ood}}\geq 70\degree$, is quite large and our method outperforms the baselines consistently.
@@
Table F.1: Detailed Results on Objaverse-OOD. We report the separate OOD evaluation results of the two sets of experiments with input views’ maximum elevations $\phi_{\textrm{max}}=(10\degree,20\degree)$. The average results are used in Tab. 1. Colors indicate the 1st, 2nd, and 3rd best-performing models.
@@
Nerfbusters (Warburg et al., 2023) 15.25 0.643 0.318 18.49 0.735 0.255 16.87 0.689 0.287
SyncDreamer (Liu et al., 2023b) 12.47 0.542 0.385 12.47 0.542 0.382 12.47 0.542 0.384
EscherNet (Kong et al., 2024) 15.99 0.606 0.267 16.20 0.611 0.258 16.09 0.608 0.263
@@
Table F.2: Detailed Results on GSO-OOD. We report the separate OOD evaluation results of the two sets of experiments with input views’ maximum elevations $\phi_{\textrm{max}}=(10\degree,20\degree)$. The average results are used in Tab. 5. Colors indicate the 1st, 2nd best-performing models.
@@
Visual Comparisons. We show more visual comparisons with all the baselines in Fig. H.2 (ShapeNet-OOD), Fig. H.4 (Objaverse-OOD), Fig. H.5 (GSO-OOD), and Fig. H.6 (Real-world OOD). We also provide a supplementary video to show the comparisons.
@@
For prior-enhanced NVS methods, SSDNeRF (Chen et al., 2023), which is designed to learn category-specific object priors in its original paper, struggles to learn cross-category priors in our training set, resulting in severe artifacts. Nerfbusters (Warburg et al., 2023) incorrectly identifies many structures as floaters, leading to the removal of significant parts of objects. While FSGS (Zhu et al., 2024) achieves a more balanced densification of Gaussians and smoother depth regularization, it only offers marginal visual improvements over 3DGS and continually fails in certain scenarios. InstantSplat (Fan et al., 2024a) also cannot mitigate the artifacts caused by overfitting by using a dense point cloud initialization.
@@
Among learning-based feed-forward methods, LaRa (Chen et al., 2024a) can infer plausible geometry and textures from input views. However, its reliance on a limited number of input views and voxel representation restricts its ability to process and represent high-frequency details. Both SyncDreamer (Liu et al., 2023b) and EscherNet (Kong et al., 2024) suffer from large hallucination errors, producing results that appear visually plausible as single views but are 3D-inconsistent and misaligned with the input views.
Appendix G Baseline Implementations
MipNeRF360. We use the officially released code [^3] https://github.com/google-research/multinerf of MipNeRF360 (Müller et al., 2022) to produce its results on our OOD benchmarks. The model is trained using batch size 512 for 250k iterations, with learning rate 0.00025. The training of each scene takes approximately 7 hours on a single RTX-4090 GPU. The near and far planes used in volumetric rendering are determined by the ray bounding box intersection following the approach used in InstantNGP. The random background trick used in InstantNGP is also applied to help removing the floaters.
3D Gaussian Splatting. We use the gsplat (Ye et al., 2024) and the default hyperparameters of the toolbox. Specifically, we set the number of iterations to 30k, warm-up steps to 500, densifying and culling Gaussians every 500 step, and we stop the density control at 15k steps. For synthetic datasets, we use the ground truth camera poses provided by the Blender rendering. Given that COLMAP struggles to reconstruct certain synthetic objects due to symmetry and smooth textures, we follow the original 3D Gaussian Splatting (3DGS) approach (Kerbl et al., 2023) by randomly sampling 50,000 points within the bounding box of the objects. For our real-world iPhone captures, we first estimate camera poses using all views and then perform point triangulation, with only the input views used to estimate the initial point cloud.
InstantSplat. We use the officially released code [^6] https://github.com/NVlabs/InstantSplat of InstantSplat. Assuming known camera poses, we use the provided training and test camera extrinsics and intrinsics, keeping them fixed in both DUST3R (Wang et al., 2024) global alignment and 3DGS optimization. Each of our OOD test scene contains 32 dense training views and applying DUST3R global alignment optimization to all $32\times 31$ pairs leads to out-of-memory issues and redundant point clouds. Therefore, we use pairwise inference results only for view pairs with overlapping observations, specifically those within an interval of 16 frames. To filter redundant points, we set the depth threshold for InstantSplat’s covisibility computation to 0.05 and retain only points with prediction confidence above the 40% quantile for each image. We train 3DGS for 3,000 iterations, using random background and switching off the camera poses optimization. We find that these configurations yield optimal performance. All other settings follow the defaults in the released code.
Nerfbusters. We use the official codebase [^9] https://github.com/ethanweber/nerfbusters of Nerfbusters (Warburg et al., 2023). We first train Nerfacto (Tancik et al., 2023) for 30k steps, and then run Nerfbusters on pretrained Nerfacto models for 5k steps to remove the artifacts learned by Nerfacto. The same near and far planes calculation strategy described in MipNeRF360 is employed.
SyncDreamer. We use the official codebase [^10] https://github.com/liuyuan-pal/SyncDreamer of SyncDreamer (Liu et al., 2023b). SyncDreamer supports only single input view, which is insufficient for accurate and faithful out-of-distribution view synthesis. Note that SyncDreamer is computationally demanding—their released model is trained for four days using eight 40GB A100 GPUs, far exceeding our computational budget. For a fair comparison, we train their method on our dataset within the same GPU hours as our approach (three days with eight RTX 4090 GPUs) and use their pretrained checkpoint for initialization.
EscherNet. We use EscherNet’s released codebase [^11] https://github.com/kxhit/EscherNet. We train two models on the same ShapeNet-OOD and Objaverse-OOD training sets as our method. Their performance is then evaluated on the corresponding benchmarks. According to the original paper, training EscherNet from scratch requires six A100 GPUs for one week, which exceeds our computational budget. Similar to our approach with SyncDreamer, we initialize training with their released checkpoint pretrained on Objaverse and finetune it on our OOD datasets. During training, we set both the numbers of input and output views to three, using eight RTX 4090 GPUs with a total batch size of eight. For inference, we input all 32 input views to predict all test views at the same time. The models are finetuned for 24k steps, achieving optimal validation performance on the two OOD datasets.
LaRa. We use LaRa’s released codebase [^12] https://github.com/autonomousvision/LaRa. We train two models on the same ShapeNet-OOD and Objaverse-OOD training sets as our method and evaluate their performances on the respective benchmarks. Due to hardware limitations, LaRa can only process a maximum of four input views on RTX-4090 GPUs. To maximize scene coverage, we randomly select the four most widely separated input views. For training, we uniformly sample from all available views, incorporating both in-distribution and out-of-distribution views for the 2D image loss. During inference, we feed the model four input views and rasterize the predicted Gaussian primitives to the OOD test views. While it is possible to divide the input views into groups of four, and then combine the processed resulting primitives for rasterization, this approach compromises global consistency and yields worse performance compared to using just four views directly. Following LaRa’s default setup (Chen et al., 2024a), we train each model for 150k iterations using 8 RTX 4090 GPUs, which takes approximately 2 days. We initialize training with LaRa’s pretrained weights, which speeds up convergence and improves performance.
Other Baselines. We acknowledge that some related approaches are not included in this paper due to redundancy or infeasibility. DNGaussian (Li et al., 2024a) supervises 3DGS training with depth maps predicted by monocular depth estimators. The idea is similar to its concurrent work FSGS (Zhu et al., 2024) which, in our OOD-NVS dense capture setup, does not outperform 3DGS. CAT3D (Gao et al., 2024) and ReconFusion (Wu et al., 2024a) use internal image diffusion models for initialization but neither has released code and models. GeNVS (Chan et al., 2023) and 3DiM (Watson et al., 2023) are also publicly unavailable. To compare with methods that utilize pretrained 2D diffusion models, we evaluate SyncDreamer (Liu et al., 2023b) and EscherNet (Kong et al., 2024) which are designed to achieve better multi-view consistency. Other similar approaches like ZeroNVS (Sargent et al., 2024), Zero123 (Liu et al., 2023a), and Vivid123 (Kwak et al., 2024), SV3D (Xie et al., 2024b) only support single-view input. We also tried ViewDiff (Höllein et al., 2024), a similar 2D-prior method where each model is trained for a single object category. However, we found that ViewDiff could not converge when trained on multiple object categories in our cases. Lastly, in the category of learning-based 2D-to-3D methods, SpaRP (Xu et al., 2024), MVSplat (Chen et al., 2024b), and LatentSplat (Wewer et al., 2024) are constrained by the memory limitations of RTX 4090 GPUs, allowing a maximum of four input views. We found that training these methods on our dataset with four large-baseline input views, as we did with LaRa, resulted in failure. This is because these methods rely on overlapping input views to compute cross-view correspondences. We also attempted to partition the consecutive input views into groups and aggregate the predicted Gaussian splats after running each group independently. However, this approach introduced significant global inconsistencies. Consequently, we do not report their results due to their incompatibility with our OOD-NVS task.